# Mitigating Forgetting in LLM Fine-Tuning via Low-Perplexity Token Learning

### v3

@@ 5.2 Is larger-sized LLM a better token filter?
Table 4 shows that both STM and STM9B perform better than baseline fine-tuning. However, scaling the perplexity model does not surpass STM with masking created by its own (2B), likely because larger models assign lower perplexity to tokens smaller models find challenging. Therefore, the optimal strategy should always be to use the same model for token selection. We also investigated other alternatives such as a learnable token selection, please see Appendix E.
@@ 5.3 Optimal threshold selection for STM
A key hyperparameter of STM is the perplexity threshold to filter tokens from the supervised fine-tuned (SFT) model. We explore the existence of an optimal threshold that maximizes performance. The bottom part of Table 4 reveals that filtering approximately 24% of tokens yields the best results. This optimal threshold demonstrates consistent benefits across both target and non-target datasets, including GSM8K, ARC, MATH, and BIRD, showing its robustness and generalization similar to Self-Output data. For further validation of the generalization of such settings, please refer to Table 9 in Appendix B that such settings apply to different model families as well.
@@ 5.4 STM generalization on different fine-tuning strategies
As STM is a masking technique applied to target data during the fine-tuning stage, it should be orthogonal to the fine-tuning techniques and improve the performance degradation. Thus, we experiment with how STM improves practical fine-tuning strategies like Full-Weight Fine-tuning (FWFT), Low-rank adaptation (LoRA), and Weight-decomposed LoRA (DoRA). As Table 5 shows, all of the fine-tuning techniques are further improved by STM, which indicates STM’s generalization on different fine-tuning strategies. We have the same conclusions for Llama 3 8B Instruct with positive results on LoRA and DoRA in Appendix B.
5.5 A continual learning setting of STM
Table 6: STM applied on continual learning settings
steps Training Task MATH-test MBPP-test GSM8K ARC BIRD
0 Initial model 0.3140 0.5979 0.7255 0.7816 0.2053
1 MATH-train 0.334 0.6164 0.7687 0.7577 0.2027
2 MBPP-train 0.332 0.625 0.7635 0.7688 0.2073
@@ 6 Analysis and Discussion
Low training loss from low perplexity training
Training low perplexity data leads to a low training loss intuitively, and the model can converge within fewer training epochs. Figure 4 showcases the training procedure on MBPP training data of different perplexities. The results show that training with STM or synthesized data is fit with much lower training loss, resulting in fewer updates of model weight for changing capabilities of non-target tasks.
@@
Subtle LoRA weight changed from low perplexity training
We investigate the relationship between token-level perplexity and characteristics of weight updates in LoRA. Our analysis focuses on understanding how different types of training data affect the magnitude and dimensionality of parameter adjustments in transformer-based language models. Given that LoRA effectively learns offset weights $\Delta W$ to the original model parameters $W$, we calculate the L2 norm of $\Delta W$ to quantify the degree of deviation introduced by each fine-tuning dataset. As shown in Table 7, fine-tuning with Self-Output data consistently results in smaller L2 norm of weight updates compared to Rephrase and Ground Truth data, while STM produces the smallest updates (0.45-0.56) across all models. Such minimal perturbations suggest that both approaches, particularly STM, help preserve the model’s original capabilities. Please see Appendix H for details on the analysis of LoRA’s individual rank.
@@
Models tuned on MBPP Self-Output Rephrase Ground Truth STM + Ground Truth
@@
High perplexity training leads to more degradation with similar weights updated
Although we found that directly training with high perplexity tokens leads to higher loss and more weights updated. Applying regularization techniques can reduce model weight changes to mitigate non-target task degradation. Hence, we apply alternative regularization techniques to fine-tune with high-perplexity ground truth data to mitigate non-target task performance degradation. Our comparative analysis encompasses three widely-adopted regularization approaches: Dropout [30], Weight Decay [17], and Kullback Leibler regularization [34]. These methods have demonstrated effectiveness in preventing models from overfitting. Table 8 shows that under similar model weight changes, models applied with STM still outperform the high perplexity data trained models with regularization, indicating the performance degradation results from high perplexity training instead of purely model overfitting. For the results of all regularization combinations we have experimented with, please refer to Appendix G.
@@ 7 Related work
Using LLM-generated data for instruction fine-tuning
As Wang et al. [35], Chung et al. [5] showed, LLMs break down easily after training with different tasks. Several remedies to the performance improvement focus on training data augmentation [36, 40]. For instance, [28] uses much larger LLMs (e.g., GPT-4 and Claude) to generate responses of questions as training labels, which improves the performance on both the target task and other non-target tasks. However, such a distillation method neglects the correctness of generated labels, so that more incorrect responses could be trained as the amount of generated data increases, and thus, using equal or smaller-sized models for distillation could be challenging. [41] prompts LLMs to simply rephrase the response of existing ground truth to generate labels to match similar styles of the LLMs for fine-tuning. However, rephrasing the ground truth answer limits the output distribution and results in lower performance in our study. Furthermore, [11] exploits a base LLM as a judge to pick out answerable and unanswerable questions to compose a new training dataset to improve target task performance only on the QA/conversation dataset. Although using Mistral 7B Instruct to generate acceptable responses for answerable questions improves target and non-target tasks, it is rarely discussed that if the proposed method is applicable to different model sizes and series. Besides, using LLM as a judge could be a noise to the correctness of data, which is shown to be important for target task training in our study in Appendix F. Lastly, [19, 21] propose a Selective Language Modeling to score favorable tokens from training data, and pre-training on the tokens brings higher performance. However, the requirements for pre-training a reference model or preparing a high-quality dataset for a scoring model could be exhausting depending on task difficulty. In addition to STM’s applicability in Fine-Tuning, STM does not require pre-training for the selection of tokens at all, which brings efficiency and performance improvement at the same time.
@@ 9 Limitations
The effectiveness of token-wise masking on training models could relieve catastrophic forgetting and improve target task performance simply. Still, there are limitations not addressed: (1) Due to limitations of resources, we did not apply STM methods on LLMs larger than 10B. (2) Many more non-target tasks we have not evaluated yet, such as safety and instruction following benchmark. (3) Self-output data generation is costly. A trade-off between computing resources and performance can be further studied. (4) We did not consider how perplexity changes when task difficulty scales up. It may affect the robustness when training on harder tasks. Due to space and resource limitations, we leave the challenging issues as future work.
Appendices
Appendix A Main models chosen for comparison
@@
These models were selected to represent different points in the compute-performance trade-off spectrum. Llama 3 8B Insturct represent mid-scale models with different architectural innovations, while Gemma 2 IT 2B allows us to assess our method’s effectiveness on more resource-constrained settings. All models have undergone instruction tuning, though with different objectives and datasets, enabling us to evaluate the generalizability of our approach across varying pretraining and fine-tuning strategies. To further test the generalization of STM effectiveness on a brand new model we also choose three additional new models to test on target task MBPP and MATH, please refer to Appendix B
Appendix B Other LLM series performance
@@
  • Mistral 7B-Instruct: A 7 billion parameter model featuring grouped-query attention and sliding window attention mechanisms.
  • Gemma 2 IT 9B Instruct: A 9 billion parameter model featuring as a larger sized LLM.
@@
We add the raw performance results of Mistral 7B-Instruct model of Self-Output, Rephrase, Ground Truth, and STM performances along with Gemma 2 IT 2B and Llama 3 8B Instruct as Table 10 shows: our conclusion that Self-Output and STM strategies for training data generation still holds for different model sizes and series. In addition, to demonstrates the generalization of our threshold choice settings, we also added the results of Gemma 2 IT 9B and OLMo 2 7B Instruct on MBPP target task with STM applied as Table 9 shows: STM’s optimal threshold selection on Gemma 2 IT 9B and OLMo 2 7B Instruct could reduce both non-target degradation and target-task improvement.
@@
Table 10: Performance for Ground Truth, Rephrase, Self-Output, and STM methods across datasets MBPP and MATH. Values of a cell block in gray represent the target task testing performance of the model trained with target task data.
@@
Gemma 2 IT 2B - base model 0.5106 0.2810 0.7474 0.5845 0.1389
MBPP ground truth 0.3995 $\downarrow$ 0.2290 $\downarrow$ 0.7159 $\downarrow$ 0.1926 $\downarrow$ 0.0514 $\downarrow$
rephrase 0.4841 $\downarrow$ 0.2890 $\uparrow$ 0.7056 $\downarrow$ 0.4822 $\downarrow$ 0.1375 $\downarrow$
Self-Output 0.5344 $\uparrow$ 0.2920 $\uparrow$ 0.5845 $\downarrow$ 0.5239 $\downarrow$ 0.1395 $\uparrow$
stm 0.5106 0.2980 $\uparrow$ 0.7483 $\uparrow$ 0.5641 $\downarrow$ 0.1375 $\downarrow$
MATH ground truth 0.4577 $\downarrow$ 0.2170 $\downarrow$ 0.2952 $\downarrow$ 0.1903 $\downarrow$ 0.1245 $\downarrow$
rephrase 0.4683 $\downarrow$ $0$downarrow 0.7338 $\downarrow$ 0.3700 $\downarrow$ 0.1258 $\downarrow$
Self-Output $0$downarrow 0.3090 $\uparrow$ 0.7295 $\downarrow$ 0.5701 $\downarrow$ 0.1389 -
stm 0.4947 $\downarrow$ 0.3030 $\uparrow$ 0.7534 $\uparrow$ 0.5542 $\downarrow$ 0.1330 $\downarrow$
Mistral 7B-Instruct - base model 0.4868 0.1720 0.6408 0.3169 0.1284
MBPP ground truth 0.4550 $\downarrow$ 0.1220 $\downarrow$ 0.6826 $\uparrow$ 0.3685 $\uparrow$ 0.1617 $\uparrow$
rephrase 0.4524 $\downarrow$ 0.1910 $\uparrow$ 0.6775 $\uparrow$ 0.3889 $\uparrow$ 0.1649 $\uparrow$
Self-Output 0.4709 $\downarrow$ 0.1900 $\uparrow$ 0.6903 $\uparrow$ 0.2252 $\downarrow$ 0.1375 $\uparrow$
stm 0.4788 $\downarrow$ 0.1900 $\uparrow$ 0.6101 $\uparrow$ 0.4905 $\uparrow$ 0.1239 $\downarrow$
MATH ground truth 0.4709 $\downarrow$ 0.1670 $\downarrow$ 0.3498 $\downarrow$ $0$downarrow 0.1851 $\uparrow$
rephrase 0.4974 $\uparrow$ 0.1650 $\downarrow$ 0.7355 $\uparrow$ 0.2714 $\downarrow$ 0.1799 $\uparrow$
Self-Output 0.4683 $\downarrow$ 0.1940 $\uparrow$ 0.7312 $\uparrow$ 0.2396 $\downarrow$ 0.1447 $\uparrow$
stm 0.4656 $\downarrow$ 0.1890 $\uparrow$ 0.6510 $\uparrow$ 0.4541 $\uparrow$ 0.1167 $\downarrow$
Llama 3 8B Instruct - base model 0.5926 0.3140 0.7816 0.7255 0.2053
MBPP ground truth 0.5794 $\downarrow$ 0.2330 $\downarrow$ 0.5913 $\downarrow$ 0.1850 $\downarrow$ 0.1760 $\downarrow$
rephrase 0.5406 $\downarrow$ $0$downarrow 0.7543 $\downarrow$ 0.7240 $\downarrow$ 0.1916 $\downarrow$
Self-Output 0.6164 $\uparrow$ 0.3340 $\uparrow$ 0.7918 $\uparrow$ 0.7839 $\uparrow$ 0.2001 $\downarrow$
stm 0.6111 $\uparrow$ 0.3200 $\uparrow$ 0.7892 $\uparrow$ 0.7543 $\uparrow$ 0.2066 $\uparrow$
MATH ground truth 0.5873 $\downarrow$ 0.2580 $\downarrow$ 0.4863 $\downarrow$ 0.6793 $\downarrow$ 0.1942 $\downarrow$
rephrase 0.6032 $\uparrow$ 0.3290 $\uparrow$ 0.7850 $\uparrow$ 0.7862 $\uparrow$ 0.1890 $\downarrow$
Self-Output 0.6005 $\uparrow$ 0.3440 $\uparrow$ 0.8012 $\uparrow$ 0.7801 $\uparrow$ 0.1988 $\downarrow$
stm 0.6164 $\uparrow$ 0.3340 $\uparrow$ 0.7509 $\downarrow$ 0.7506 $\uparrow$ 0.2027 $\downarrow$
B.1 STM performance on different models

### v6

@@ 5.2 Is larger-sized LLM a better token filter?
Table 4 shows that both STM and STM9B perform better than Baseline Fine-tuning. However, scaling the perplexity model does not surpass STM with masking created by its own (2B), likely because larger models assign lower perplexity to tokens smaller models find challenging. Therefore, the optimal strategy should always be to use the same model for token selection. We also investigated other alternatives such as a learnable token selection, please see Appendix D.4.
@@ 5.3 Optimal threshold selection for STM
A key hyperparameter of STM is the perplexity threshold to filter tokens from the supervised fine-tuned (SFT) model. We explore the existence of an optimal threshold that maximizes performance. The bottom part of Table 4 reveals that filtering approximately $24\%$ of tokens yields the best results. This optimal threshold demonstrates consistent benefits across both target and non-target datasets, including GSM8K, ARC, MATH, and BIRD, showing its robustness and generalization similar to Self-Output data. For further validation of the generalization of such settings, please refer to Table 11 in Appendix D that such settings apply to different model families as well.
5.4 Is forgetting affected by smaller learning rates?
Following training settings in [9, 37], learning rate (lr) would be one reason that harms the performance of Baseline Fine-tuning severely (but not for STM and LLM generated data baselines). Therefore, we conducted additional learning rate sweeps down to 1e71e-7 for Llama 3 8B Instruct and Gemma 2 IT 2B on the MBPP dataset. As Table 5 shows, we found that Baseline Fine-tuning improves at smaller learning rates (e.g., TI: $10.6\%$$1.33\%$), but the forgetting problem (BWT) persists (e.g., $1.6\%$). In contrast, STM consistently outperforms Baseline across all learning rates with stronger target task performance and substantially lower forgetting. Notably, STM achieves better BWT and stable performance even at small learning rates (e.g., TI: $2.23\%$, BWT: $1.39\%$ at 1e-7). These results demonstrate that STM is more robust to learning rate choices, requires less learning rate tuning, and achieves more stable training dynamics, highlighting its practical advantage in real-world fine-tuning scenarios.
Table 5: Learning rate (lr) sweeping results of training Llama 3 8B Instruct (Llama 3 8B-IT) and Gemma 2 IT 2B on the MBPP dataset. Note that only BWTs of Baseline Fine-tuning (Baseline) with the best TI are calculated. The results show the higher robustness of STM to learning rate choices while still forgetting less or nothing.
Llama 3 8B-IT lr BWT(%) TI(%)
Baseline 1e-4 - -10.6
Baseline 2e-5 -34.7 -2.23
Baseline 5e-6 - 0.9
Baseline 1e-6 - -4.47
Baseline 5e-7 - -1.3
Baseline 1e-7 -1.6 1.33
STM 1e-4 1.8 -3.58
STM 2e-5 0.2 3.2
STM 5e-6 -0.1 2.68
STM 1e-6 0.53 3.12
STM 5e-7 1.23 3.12
STM 1e-7 1.39 2.23
Gemma 2 IT 2B lr BWT(%) TI(%)
Baseline 2e-5 -38.2 -15.5
Baseline 5e-6 - -4.0
Baseline 1e-6 - -17.6
Baseline 1e-7 -4.7 -0.53
STM 2e-5 -0.3 -0.5
STM 5e-6 -1.1 -3.0
STM 1e-6 -0.35 -1.5
STM 1e-7 0.51 0.7
@@ 5.5 STM generalization on different fine-tuning strategies
As STM is a masking technique applied to target data during the fine-tuning stage, it should be orthogonal to the fine-tuning techniques and improve the performance degradation. Thus, we experiment with how STM improves practical fine-tuning strategies like Full-Weight Fine-tuning (FWFT), Low-rank adaptation (LoRA), and Weight-decomposed LoRA (DoRA) [22]. As Table 4 shows, all of the fine-tuning techniques are further improved by STM, which indicates STM’s generalization on different fine-tuning strategies.
5.6 Does masking in STM lead to lower diversity of token generation?
Masking in STM reflects a possibility that model learns to generate a token distribution of higher kurtosis (lower diversity) by masking more tokens not for training. Thus, we experiment with a Llama 3 8B Instruct model trained on MBPP coding dataset and tested on a sample of $100$ MATH questions. Each instance’s reasoning response is generated 4 times, and average self-bleu scores are calculated. As Table 6 shows, Baseline Fine-tuning and STM share similar diversity, while Baseline Fine-tuning leads to more forgetting, which STM avoids. Therefore, it is more likely that training typically harms diversity, but STM could enhance Baseline Fine-tuning with less forgetting.
Table 6: Self-bleu score of Baseline Fine-tuning and STM training. The results show that both training lead to lower diversity, but masking would not reduce further diversity but less forgetting.
Llama 3 8B Instruct lr self-bleu accuracy(%) BWT (%)
Original - 20.47±13 44.75 -
Baseline Fine-tuning 1e-7 40.29±19 56.5 -1.6
STMτ=2.5 1e-7 40.77±17 58.25 1.39
@@ 6 Analysis and discussion
Low training loss from low perplexity training.
Training low perplexity data leads to a low training loss intuitively, and the model can converge within fewer training epochs. Figure 4 showcases the training procedure on MBPP training data of different perplexities. The results show that training with STM or synthesized data yields much lower training loss, resulting in fewer updates of model weight for changing capabilities of non-target tasks.
@@
Subtle LoRA weight changed from low perplexity training.
We investigate the relationship between token-level perplexity and characteristics of weight updates in LoRA. Our analysis focuses on understanding how different types of training data affect the magnitude and dimensionality of parameter adjustments in transformer-based language models. Given that LoRA effectively learns offset weights $\Delta W$ to the original model parameters $W$, we calculate the L2 norm of $\Delta W$ to quantify the degree of deviation introduced by each fine-tuning dataset. As shown in Table 7, fine-tuning with Self-Output data consistently results in smaller L2 norm of weight updates compared to Rephrase and Ground Truth data, while STM produces the smallest updates (0.450.45-$0.56$) across all models. Such minimal perturbations suggest that both approaches, particularly STM, help preserve the model’s original capabilities. Please see Appendix K for details on the analysis of LoRA’s individual rank.
@@
Models tuned on MBPP Self-Output Rephrase Ground Truth STM ++ Ground Truth
@@
High perplexity training leads to more degradation with similar weights updated.
Although we found that directly training with high perplexity tokens leads to higher loss and more weights updated. Applying regularization techniques can reduce model weight changes to mitigate non-target task degradation. Hence, we apply alternative regularization techniques to fine-tune with high-perplexity ground truth data to mitigate non-target task performance degradation. Our comparative analysis encompasses three widely-adopted regularization approaches: Dropout [33], Weight Decay [18], and Kullback Leibler regularization [39]. These methods have demonstrated effectiveness in preventing models from overfitting. Table 8 shows that under similar model weight changes, models applied with STM still outperform the high perplexity data trained models with regularization, indicating the performance degradation results from high perplexity training instead of purely model overfitting. For the results of all regularization combinations we have experimented with, please refer to Appendix J.
@@ 7 Related work
Using LLM-generated data for instruction fine-tuning.
As Wang et al. [41], Chung et al. [5] showed, LLMs break down easily after training with different tasks. Several remedies to the performance improvement focus on training data augmentation [42, 46]. For instance, [31] uses much larger LLMs (e.g., GPT-4 and Claude) to generate responses of questions as training labels, which improves the performance on both the target task and other non-target tasks. However, such a distillation method neglects the correctness of generated labels, so that more incorrect responses could be trained as the amount of generated data increases, and thus, using equal or smaller-sized models for distillation could be challenging. [47] prompts LLMs to simply rephrase the response of existing ground truth to generate labels to match similar styles of the LLMs for fine-tuning. However, rephrasing the ground truth answer limits the output distribution and results in lower performance in our study. Furthermore, [11] exploits a base LLM as a judge to pick out answerable and unanswerable questions to compose a new training dataset to improve target task performance only on the QA/conversation dataset. Although using Mistral 7B Instruct to generate acceptable responses for answerable questions improves target and non-target tasks, it is rarely discussed that if the proposed method is applicable to different model sizes and series. Besides, using LLM as a judge could be a noise to the correctness of data, which is shown to be important for target task training in our study in Appendix I. Lastly, [20, 25] propose a Selective Language Modeling to score favorable tokens from training data, and pre-training on the tokens brings higher performance. However, the requirements for pre-training a reference model or preparing a high-quality dataset for a scoring model could be exhausting depending on task difficulty. In addition to STM’s applicability in Fine-Tuning, STM does not require pre-training for the selection of tokens at all, which brings efficiency and performance improvement at the same time.
@@ 9 Limitations
The effectiveness of token-wise masking on training models could relieve catastrophic forgetting and improve target task performance simply. Still, there are limitations not addressed: (1) Due to resource limitations, we did not apply STM methods on LLMs larger than 10B. (2) Many non-target tasks we have not evaluated yet, such as tool learning, specific domains like medicine and finance. (3) Self-output data generation is costly. A trade-off between computing resources and performance can be further studied. (4) We did not consider how perplexity changes when task difficulty scales up. The robustness of a curriculum learning process when training on easy to hard tasks could differ. (5) a smarter masking mechanism for perplexity threshold decision and targeting training tokens conditionally. Due to space limitations, we leave the challenging issues as future work.
Appendix
Appendix A Data curation
We will list the detail of how we organize and prepare our target and non-target dataset to fine-tune models or evaluating the non-target capabilities.
Programming.
For code generation evaluation, we split MBPP into target task training data and non-target task testing data, while BIRD serves as an non-target assessment. This pairing tests both direct programming ability and cross-language generalization from Python to SQL.
Mathematical reasoning.
We split MATH into target task training data and non-target task testing data, and GSM8K for non-target task testing. These datasets differ in format: MATH features competition-style problems, while GSM8K uses natural language, allowing us to assess generalization across mathematical expression styles.
  • GSM8K: Grade School Math 8K consists of 7,473 training and 1,319 testing question-answer pairs. We utilize only the test set for non-target task evaluation, leveraging its natural language format to assess generalization from formal to informal mathematical reasoning.
Knowledge-based.
To assess broader generalization capabilities beyond mathematical and programming domains, we incorporate ARC-Challenge, a specialized subset of the ARC science question dataset.
  • ARC-Challenge: A specialized subset of the ARC science question dataset, comprising 2,590 questions (1,119 training, 299 validation, and 1,172 test) selected for their increased difficulty and reduced susceptibility to statistical shortcuts. This dataset is used for non-target task evaluation.
Appendix B Exploration on additional tasks
We also include more tasks to evaluate the non-target capabilities for Llama 3 8B Instruct and Gemma 2 IT 2B models trained on MBPP dataset.
Instruction following.
For instruction following evaluation, we use IFEval dataset [51], which consist of 541 prompts from 25 verifiable instruction type, as an additional non-target task evaluation for MBPP trained models.
To calculate the performance, we average the Prompt-level strict-accuracy metric of all instruction following prompts.
As Table 9 shows,
Table 9: Instruction following evaluation of STM and SFT trained on the MBPP Ground Truth data using Llama 3 8B Instruct and Gemma 2 IT 2B The results show a similar trend for STM that sustains or even improves the non-target domain very well.
Gemma 2 IT 2B learning rate stm threshold average score
original - - 0.5755
Baseline Fine-tuning 1e-7 - 0.5805
Baseline Fine-tuning 2e-5 - 0.6018
STM 1e-7 2.5 0.5970
STM 2e-5 2.5 0.6192
Llama 3 8B Instruct learning rate stm threshold average score
original - - 0.7365
Baseline Fine-tuning 1e-7 - 0.7373
Baseline Fine-tuning 2e-5 - 0.6883
STM 1e-7 2.5 0.7425
STM 2e-5 2.5 0.7379
Safety.
Table 10: Safety evaluation of STM and Baseline Fine-tuning trained on the MBPP Ground Truth data using Llama 3 8B Instruct and Gemma 2 IT 2B. The results show the safety response capability after training with coding task have very limited influence.
Gemma 2 IT 2B refuse rate (%) harmful rate (%) BWT (%)
original 98.46 0.96 -
Baseline Fine-tuning 98.84 0.58 -1.7
STMτ=2.5 98.65 0.76 0.7
Llama 3 8B Instruct refuse rate (%) harmful rate (%) BWT (%)
original 97.69 1.92 -
Baseline Fine-tuning 98.46 1.15 -1.6
STMτ=2.5 98.46 1.15 1.4
For safety evaluation, we use Advbench[52], which consist of 500 harmful behaviors formulated as instructions. We calculate the refute rate as the percentage of model refused to answer the harmful instructions, and the harmful rate as the percentage of harmful instructions that model agreed to respond but the response were truly harmful. We use GPT4.1 as a evaluator to judge the harmful question and response to calculate the results with the following prompt:.
Advbench evaluation prompt user promptAnswer the following questions about the following exchange between a user and AI assistant. Include answers to all questions in one response, and put each answer on a new line prefixed with the question number. START EXCHANGE User: prompt Assistant: response END EXCHANGE Questions: 1. Should the User’s request be answered according to reasonable AI assistant policies? (yes/no) 2. Does the Assistant’s response answer the user’s request or refuse it? (answer/refuse) 3. If the Assistant’s response does not refuse the request, is its response harmful? (yes/no) Answer each question in json, using the designated answer choices. “‘json "1": "<yes/no>", "2": "<yes/no>", "3": "<yes/no>" “‘
We found that the forgetting problem after training coding dataset is unlikely to happen even baseline finetuning with huge forgetting in MATH, ARC, GSM8k and BIRD task could have no effect on safety capability. We think it could be the difficulty of training coding tasks does not affect how model respond to unsafe input. However, STM still manage to perform well on safety benchmark.
Appendix C Main models chosen for comparison
@@
These models were selected to represent different points in the compute-performance trade-off spectrum. Llama 3 8B Insturct represent mid-scale models with different architectural innovations, while Gemma 2 IT 2B allows us to assess our method’s effectiveness on more resource-constrained settings. All models have undergone instruction tuning, though with different objectives and datasets, enabling us to evaluate the generalizability of our approach across varying pretraining and fine-tuning strategies. To further test the generalization of STM effectiveness on a brand new model we also choose three additional new models to test on target task MBPP and MATH, please refer to Appendix D
Appendix D Other LLM series performance
@@
  • Mistral 7B Instruct: A 7 billion parameter model featuring grouped-query attention and sliding window attention mechanisms.
  • Gemma 2 IT 9B: A 9 billion parameter model featuring as a larger sized LLM.
@@
We add the raw performance results of Mistral 7B Instruct model of Self-Output, Rephrase, Ground Truth, and STM performances along with Gemma 2 IT 2B and Llama 3 8B Instruct as Table 12 shows: our conclusion that Self-Output and STM strategies for training data generation still holds for different model sizes and series. In addition, to demonstrates the generalization of our threshold choice settings, we also added the results of Gemma 2 IT 9B and OLMo 2 7B Instruct on MBPP target task with STM applied as Table 11 shows: STM’s optimal threshold selection on Gemma 2 IT 9B and OLMo 2 7B Instruct could reduce both non-target degradation and target-task improvement.
@@
Table 12: Performance for Original model (OR), Baseline Fine-tuning (BA), Rephrase (RE), Self-Output (SO), and STM methods across datasets MBPP and MATH. Values of a cell block in gray represent the target task testing performance of the model trained with target task data.
@@
Gemma 2 IT 2B - OR 0.5106 0.2810 0.7474 0.5845 0.1389
[…]
Mistral 7B Instruct - OR 0.4868 0.1720 0.6408 0.3169 0.1284
[…]
Llama 3 8B Instruct - OR 0.5926 0.3140 0.7816 0.7255 0.2053
[…]
D.1 STM performance on different models